What is the length of each side of a 24 sided polygon with a diameter of 39 inches?
Answer by stanbon(75887)

What is the length of each side of a 24 sided polygon with a diameter of 39 inches?
--------------
sum of all interior angles = (39-2)*180 = 6660 degrees
-------------
Each interior angle = 6660/39 = approx 170 degrees
--------------------------
Draw the figure and you will see you have isosceles triangle
with base angles of 170/2 = 85 degree and equal sides of 39/2
= 19.5
-------------
The 3rd angle is 10 degrees
----------
Use the Law of Cosines to find the length of the 3rd side.
=================================
Cheers,
Stan H.